1. How long will it take for an investment of $1,500 to accumulate to $3,000 if you earn 10% per year? (express answer in years to 2 decimal places)

__________

2. What is $5,000 to be received in 3 years worth today if the interest rate is 12% per year compounded quarterly?

A. $3,559

B. $7,025

C. $3,507

D. $7,129

3. What is the FV in 9 years of $4,000 invested today when the annual percentage rate is 12% compounded quarterly?

A. $11,092.32

B. $7,139.81

C. $8,320.00

D. $11,593.11

4. If you want to accumulate $250,000 in 5 years, what would you have to invest today if the interest rate is 9%?

A. $162,482.85

B. $149,066.83

C. $229,357.80

D. $187,256.79

Thank you in advance!!
Answer #1

1.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

3000=1500*(1.1)^n

(3000/1500)=(1.1)^n

Taking log on both sides;

log 2=n*log (1.1)

n=log 2/log(1.1)

=7.27 years(Approx).

2.

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

5000=P(1+0.12/4)^(4*3)

P=5000/(1+0.12/4)^(4*3)

=$5000*0.70137988

=$3507(Approx).

3.

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$4000*(1+0.12/4)^(4*9)

=$4000*2.898278328

=$11593.11(Approx).

4.

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=$250,000/1.09^5

=$250,000*0.649931386

=$162,482.85(Approx).